### 05 Mar 2024 v2

@@ 3.2.2 Training with Inconsistent Labels in Multimodal Fusion
Addressing label inconsistency is crucial in multimodal classification. In the PE dataset, each CT scan slice has its own label. For instance, a CT volume of size 256$\times$256$\times$100 yields 100 individual slice labels. However, the corresponding EHR data, is assigned only one single label, either positive or negative. Previous study Huang et al. (2020) established a benchmark for classifying PE dataset image data, dividing the CT volume into multiple 24 consecutive slices. If any 4 slices within any 24 consecutive slices are identified as positive, the entire volume is labeled positive. This approach does not completely solve the label inconsistency problem when fusing EHR and CT volume data, as the EHR label might not align with the image’s. Refer to Appendix F for details.
@@ 3.2.3 TNF-Based Models for PE Classification
where $\mathcal{T}$ represents the class embedding and Transformer blocks. Refer to Appendix E for detailed network structure.
@@ 4.3.1 PE Dataset
The results for the PE dataset are detailed in Table 2, with the ROC curve displayed in Fig. 4. We compared TNF-based models with single modal, multimodal fusion, and ensemble approaches. We tried two fusion models: 1) MMTM fusion, whose structure is equivalent to the model in Fig. 2 (c) while only outputs fusion result. 2) Clip fusion, based on the research of Hager et al. (2023) (refer to Appendix A for details). The TNF approach not only surpassed single-modal classification (using either image or tabular) but also outperformed multimodal fusion and ensemble in most metrics. Notably, the fusion approach demonstrated lower accuracy compared to the ensemble method, indicating that conventional multimodal fusion may be ineffective for some datasets.
@@ 6 Conclusion
Details of CLIP Fusion in the “Experiments” Section
We employed a multimodal fusion approach shown in Fig. 1 inspired by the work of Hager et al. Hager et al. (2023), which itself is based on the principles derived from CLIP radford2021learning. We employed CLIP loss for fine-tuning our pretrained models, leveraging CLIP’s exceptional performance not only in pre-training phases but also in fine-tuning scenarios, as highlighted by Dong et al. dong2022clip. The training and inference processes are outlined as follows:
Experimental results demonstrated in Table 6 indicate that single modality and ensemble surpassed the performance of CLIP fusion. The fact suggests that the efficacy of fusion methods is dependent on the datasets used. Some datasets contain multimodal data with high inter-modality correlation, which enhances the performance of fusion-based approaches. Conversely, in datasets where the multimodal data are relatively independent, fusion methods may not yield improvements.
CLIP fusion Single modality & ensemble
Ensemble Img only Tab only Ensemble Img only Tab only
AUPRC $\uparrow$ 0.838 0.816 0.832 0.888 0.848 0.842
AUROC $\uparrow$ 0.797 0.734 0.770 0.847 0.786 0.788
ACC $\uparrow$ 0.705 0.663 0.621 0.774 0.716 0.737
MCC $\uparrow$ 0.393 0.291 0.202 0.534 0.407 0.472
Recall $\uparrow$ 0.918 0.882 0.964 0.818 0.855 0.727
Table 6: Performance comparison of models w/ and w/o CLIP fusion. CLIP fusion lead to a decline in most metrics.
Figure 1: Process of CLIP fusion is adapted from Hager et al. Hager et al. (2023). During the training phase, we fine-tune both the PENet and the Tabular Transformer by employing a combination of cross-entropy loss and CLIP loss. For the inference phase, we utilize the fine-tuned PENet and Tabular Transformer to perform predictions.
A SHAP-based Attribute Selection
Figure 2: Top 250 influential attributes of the PE dataset’s EHR tabular data.
Figure 3: Closer examination of the top 20 attributes, along with their levels of contribution of the PE dataset’s EHR tabular data.
Attributes 1503 1000 500 250 100
AUROC 0.727 0.697 0.693 0.788 0.777
ACC 0.694 0.705 0.684 0.737 0.737
MCC 0.390 0.421 0.403 0.472 0.467
Table 7: Impact of selecting different number of tabular attributes on the PE dataset
The formula to calculate SHAP values is given by:
The results, presented in Table 7, indicate that a subset of 250 attributes yielded superior classification accuracy compared to larger or smaller subsets of attributes. Illustrations of the top 250 influential attributes are provided in Fig. 2, and a closer examination of the top 20 attributes, along with their levels of contribution, is depicted in Fig. 3.
C Preprocessing of NACC Dataset
  • Inclusion of the T1 modality.
  • Availability of high-resolution, thin-slice MRI scans.
Figure 4: We employed FreeSurfer Fischl (2012) for skull stripping, resizing, and intensity normalization of NACC dataset’s MRI volumes.
Additionally, we employed FreeSurfer Fischl (2012) for skull stripping, resizing, and intensity normalization of the MRI data, as illustrated in Fig. 4. After FreeSurfer’s processing, each voxel is with dimensions of 182$\times$218$\times$182 voxels.
D Selecting Model for PE Tabular Classification
D Details of TNF-Based Models for PE Classification
Figure 5: TNF-based networks on PE dataset.
E.1 Modified MMTM-based feature fusion
Initially, we carried out experiments utilizing TNF with a modified MMTM-based feature fusion approach. We employed MMTM to integrate features from image and tabular data, followed by the application of a linear layer to obtain the output of the fusion branch. The architecture of this model is illustrated in Fig. 5 (a).
E.2 Transformer Feature Fusion
Concatenation-Based Feature Fusion We also assessed the concatenation-based feature fusion of TNF on transformer architectures. In this approach, we processed the image CNN features from PENet through a transformer block to obtain transformer features, which were then concatenated with features from the tabular transformer. Subsequently, these combined features were processed through 8 transformer blocks to derive the output of the fusion branch, as illustrated in Fig. 5 (b).
TokenFusion and Cross-Modal Attention-Based Feature Fusion The architectures of models based on TokenFusion Wang et al. (2022) and Cross-Modal Attention Li and Li (2020) are presented in Fig. 5 (c). For fusing image and tabular features, we employed either TokenFusion or Cross-Modal Attention. These fused features were then processed through 8 transformer blocks to produce the output of the fusion branch.
E Details About Maximum Likelihood Selection and Label Masking
Single slice 24 slices
Img=1 Img=0 Img=1 Img=0
Tab=1 44,071 244,591 2,361 10,025
Tab=0 0 433,956 0 18,608
Table 8: Label inconsistency problem in the PE dataset: Each slice (image) within CT volumes is individually labeled, whereas the corresponding EHR data (tabular) for the entire CT volume has a single label, leading to situations where some image labels are 0 while the corresponding EHR data label is 1.
E.1 Label Inconsistency Problem
@@
Inconsistencies may arise between EHR labels and image labels within the PE dataset as shown in Table 8. Huang et al. Huang et al. (2020) set a classification benchmark for PE image data by segmenting the CT volume into groups, with each containing 24 consecutive slices. A volume is deemed positive if at least 4 slices in any of these groups are positive. Our study adheres to this approach.
@@
However, a positive PE volume doesn’t guarantee that all its slices are positive. Actually, according to Table 8, for all the volumes which are annoted as positive, only 15.27% slices are positive, and only 19.06% 24-slice groups are positive. This discrepancy leads to label inconsistencies when comparing positive EHR data with potentially negative image groups. We propose 1) label masking, and 2) maximum likelihood selection to tackle this problem.
E.2 Maximum Likelihood Selection
@@ E.3 Label Masking
E Additional Grad-CAM Results
@@
Figure 6: More Grad-CAM heatmaps. Yellow arrows indicate PE lesions.
@@
To demonstrate the superior performance of TNF-based models compared to others, additional Grad-CAM heatmaps are presented in Fig. 6. We compared five types of Grad-CAM heatmaps: 1) The fusion branch of TNF. The model is shown in Fig. 5 (a); 2) The image branch of TNF; 3) PENet; 4) MMTM fusion model; The model’s structure is equivalent to the model in Fig. 5 (a) and only outputs fusion result; 5) Clip fusion model. Yellow arrows in Fig. 6 indicate PE lesions. In short, the highlighted parts of the TNF-based models are closer to the lesions.

### 10 Mar 2024 v3

@@ 3.2.2 Training with Inconsistent Labels in Multimodal Fusion
Addressing label inconsistency is crucial in multimodal classification. In the PE dataset, each CT scan slice has its own label. For instance, a CT volume of size 256$\times$256$\times$100 yields 100 individual slice labels. However, the corresponding EHR data, is assigned only one single label, either positive or negative. Previous study Huang et al. (2020) established a benchmark for classifying PE dataset image data, dividing the CT volume into multiple 24 consecutive slices. If any 4 slices within any 24 consecutive slices are identified as positive, the entire volume is labeled positive. This approach does not completely solve the label inconsistency problem when fusing EHR and CT volume data, as the EHR label might not align with the image’s. Refer to Appendix A for details.
@@ 3.2.3 TNF-Based Models for PE Classification
where $\mathcal{T}$ represents the class embedding and Transformer blocks. Refer to Appendix B for detailed network structure.
@@ 4.3.1 PE Dataset
The results for the PE dataset are detailed in Table 2, with the ROC curve displayed in Fig. 4. We compared TNF-based models with single modal, multimodal fusion, and ensemble approaches. We tried two fusion models: 1) MMTM fusion, whose structure is equivalent to the model in Fig. 2 (c) while only outputs fusion result. 2) Clip fusion, based on the research of Hager et al. (2023) (refer to Appendix C for details). The TNF approach not only surpassed single-modal classification (using either image or tabular) but also outperformed multimodal fusion and ensemble in most metrics. Notably, the fusion approach demonstrated lower accuracy compared to the ensemble method, indicating that conventional multimodal fusion may be ineffective for some datasets.
@@ 6 Conclusion
A Details About Maximum Likelihood Selection and Label Masking
A.1 Label Inconsistency Problem
@@
Inconsistencies may arise between EHR labels and image labels within the PE dataset as shown in Table A. Huang et al. Huang et al. (2020) set a classification benchmark for PE image data by segmenting the CT volume into groups, with each containing 24 consecutive slices. A volume is deemed positive if at least 4 slices in any of these groups are positive. Our study adheres to this approach.
@@
However, a positive PE volume doesn’t guarantee that all its slices are positive. Actually, according to Table A, for all the volumes which are annoted as positive, only 15.27% slices are positive, and only 19.06% 24-slice groups are positive. This discrepancy leads to label inconsistencies when comparing positive EHR data with potentially negative image groups. We propose 1) label masking, and 2) maximum likelihood selection to tackle this problem.
A.2 Maximum Likelihood Selection
@@ A.3 Label Masking
Single slice 24 slices
Img=1 Img=0 Img=1 Img=0
Tab=1 44,071 244,591 2,361 10,025
Tab=0 0 433,956 0 18,608
Table A: Label inconsistency problem in the PE dataset: Each slice (image) within CT volumes is individually labeled, whereas the corresponding EHR data (tabular) for the entire CT volume has a single label, leading to situations where some image labels are 0 while the corresponding EHR data label is 1.
B Details of TNF-Based Models for PE Classification
Figure A: TNF-based networks on PE dataset.
B.1 Modified MMTM-based feature fusion
Initially, we carried out experiments utilizing TNF with a modified MMTM-based feature fusion approach. We employed MMTM to integrate features from image and tabular data, followed by the application of a linear layer to obtain the output of the fusion branch. The architecture of this model is illustrated in Fig. A (a).
B.2 Transformer Feature Fusion
Concatenation-Based Feature Fusion We also assessed the concatenation-based feature fusion of TNF on transformer architectures. In this approach, we processed the image CNN features from PENet through a transformer block to obtain transformer features, which were then concatenated with features from the tabular transformer. Subsequently, these combined features were processed through 8 transformer blocks to derive the output of the fusion branch, as illustrated in Fig. A (b).
TokenFusion and Cross-Modal Attention-Based Feature Fusion The architectures of models based on TokenFusion Wang et al. (2022) and Cross-Modal Attention Li and Li (2020) are presented in Fig. A (c). For fusing image and tabular features, we employed either TokenFusion or Cross-Modal Attention. These fused features were then processed through 8 transformer blocks to produce the output of the fusion branch.
C Details of CLIP Fusion in the “Experiments” Section
We employed a multimodal fusion approach shown in Fig. B inspired by the work of Hager et al. Hager et al. (2023), which itself is based on the principles derived from CLIP radford2021learning. We employed CLIP loss for fine-tuning our pretrained models, leveraging CLIP’s exceptional performance not only in pre-training phases but also in fine-tuning scenarios, as highlighted by Dong et al. dong2022clip. The training and inference processes are outlined as follows:
Experimental results demonstrated in Table B indicate that single modality and ensemble surpassed the performance of CLIP fusion. The fact suggests that the efficacy of fusion methods is dependent on the datasets used. Some datasets contain multimodal data with high inter-modality correlation, which enhances the performance of fusion-based approaches. Conversely, in datasets where the multimodal data are relatively independent, fusion methods may not yield improvements.
CLIP fusion Single modality & ensemble
Ensemble Img only Tab only Ensemble Img only Tab only
AUPRC $\uparrow$ 0.838 0.816 0.832 0.888 0.848 0.842
AUROC $\uparrow$ 0.797 0.734 0.770 0.847 0.786 0.788
ACC $\uparrow$ 0.705 0.663 0.621 0.774 0.716 0.737
MCC $\uparrow$ 0.393 0.291 0.202 0.534 0.407 0.472
Recall $\uparrow$ 0.918 0.882 0.964 0.818 0.855 0.727
Table B: Performance comparison of models w/ and w/o CLIP fusion. CLIP fusion lead to a decline in most metrics.
Figure B: Process of CLIP fusion is adapted from Hager et al. Hager et al. (2023). During the training phase, we fine-tune both the PENet and the Tabular Transformer by employing a combination of cross-entropy loss and CLIP loss. For the inference phase, we utilize the fine-tuned PENet and Tabular Transformer to perform predictions.
D SHAP-based Attribute Selection
Figure C: Top 250 influential attributes of the PE dataset’s EHR tabular data.
Figure D: Closer examination of the top 20 attributes, along with their levels of contribution of the PE dataset’s EHR tabular data.
Attributes 1503 1000 500 250 100
AUROC 0.727 0.697 0.693 0.788 0.777
ACC 0.694 0.705 0.684 0.737 0.737
MCC 0.390 0.421 0.403 0.472 0.467
Table C: Impact of selecting different number of tabular attributes on the PE dataset
The formula to calculate SHAP values is given by:
The results, presented in Table C, indicate that a subset of 250 attributes yielded superior classification accuracy compared to larger or smaller subsets of attributes. Illustrations of the top 250 influential attributes are provided in Fig. C, and a closer examination of the top 20 attributes, along with their levels of contribution, is depicted in Fig. D.
E Preprocessing of NACC Dataset
  • Inclusion of the T1 modality.
  • Availability of high-resolution, thin-slice MRI scans.
Figure E: We employed FreeSurfer Fischl (2012) for skull stripping, resizing, and intensity normalization of NACC dataset’s MRI volumes.
Additionally, we employed FreeSurfer Fischl (2012) for skull stripping, resizing, and intensity normalization of the MRI data, as illustrated in Fig. E. After FreeSurfer’s processing, each voxel is with dimensions of 182$\times$218$\times$182 voxels.
F Selecting Model for PE Tabular Classification
G Additional Grad-CAM Results
@@
Figure F: More Grad-CAM heatmaps. Yellow arrows indicate PE lesions.
@@
To demonstrate the superior performance of TNF-based models compared to others, additional Grad-CAM heatmaps are presented in Fig. F. We compared five types of Grad-CAM heatmaps: 1) The fusion branch of TNF. The model is shown in Fig. A (a); 2) The image branch of TNF; 3) PENet; 4) MMTM fusion model; The model’s structure is equivalent to the model in Fig. A (a) and only outputs fusion result; 5) Clip fusion model. Yellow arrows in Fig. F indicate PE lesions. In short, the highlighted parts of the TNF-based models are closer to the lesions.